Dear sir, I was facing problems in this question. Calculate the total number of oxygen atoms present in 1.80g of glucose c6h12o6. In my book the ans is 3.6*10^22 but I am getting 3.6*10^21 Pls tell me where I have gone wrong.

Dear sir,
I was facing problems in this question.
Calculate the total number of oxygen atoms present in 1.80g of glucose c6h12o6.
In my book the ans is 3.6*10^22 but I am getting 3.6*10^21
Pls tell me where I have gone wrong.

Grade:12

2 Answers

bharat bajaj IIT Delhi
9 years ago
Molecular weight of Glucose = 180One molecule of glucose = 6 atoms of oxygenTotal molecules of glucose = 1.8/180 X 6.022*10^23 = 6.022*10^21Total oxygen atoms = 6 * 6.022*10^21 = 36 * 10^21 = 3.6*10^22Hope you find your mistake.Thanks
Bharat BajajIIT Delhi
askiitians faculty
Jai Mahajan
9 years ago
Sir what if the question was to find number of oxygen molecules? Would we multiply 6.022*10^21 by 3?
Thanks a lot